OpenStudy (anonymous):

Help please! 1/2 x+1/4 y =-4 -2/3 x+1/2 y =7 What is X? and Y?

OpenStudy (texaschic101):

do you want to do this with the fractions included, or do you want to do a little extra step and turn the fractions into whole numbers where you do not have to deal with the fractions ?

OpenStudy (anonymous):

Either or If its easier keep it in fractions

OpenStudy (texaschic101):

I think it is easier to turn them to whole numbers. Let me show you how.. 1/2x + 1/4y = -4 --- multiply entire equation by common denominator of 4 (4) 1/2x + (4) 1/4y = (4) -4 2x + 4y = -16 now for the other one.. -2/3x + 1/2y = 7 --- multiply by common denominator of 6 (6) -2/3x + (6) 1/2y = (6)7 -4x + 3y = 42 now we have : 2x + 4y = -16 -4x + 3y = 42 now we can either solve by elimination or substitution ...which do you prefer...and do you have any questions so far ?

OpenStudy (texaschic101):

wait...I caught a mistake...first equation, when multiplied by 4, should be 2x + y = -16

OpenStudy (texaschic101):

so lets use substitution.. 2x + y = -16 y = -2x - 16 now sub -2x - 16 in for y in the other equation. -4x + 3y = 42 -4x + 3(-2x - 16) = 42 -4x - 6x - 48 = 42 -4x - 6x = 42 + 48 -10x = 90 x = 90/-10 x = -9 y = -2x - 16 y = -2(-9) - 16 y = 18 - 16 y = 2 I might have done something wrong...its not checking correctly...hold on

OpenStudy (texaschic101):

1/2x + 1/4y = -4 1/2(-9) + 1/4(2) = -4 -9/2 + 2/4 = -4 -18/4 + 2/4 = -4 -16/4 = -4 -4 = -4 (correct) oh...lol...nevermind...it is correct. x = -9 and y = 2

OpenStudy (anonymous):

Thank so much and thanks for showing me how its done!!